OpenStudy (anonymous):

If f(x)=ax^2+(x-b)^2 has critical point at (1,6) , determine if this critical point is maximum or minimum.

OpenStudy (anonymous):

@ikram002p

OpenStudy (iamtheone):

http://www.wolframalpha.com/input/?i=minimum+or+maximum Happy Studying!!!!!

OpenStudy (anonymous):

@IAMTHEONE i know how to solve it but i'm not sure about my answer

OpenStudy (anonymous):

@ikram002p i just want the values of a and b to compare it with my answer

OpenStudy (ikram002p):

ugh ok i was typing it but somhow i get offline :|

OpenStudy (anonymous):

ok can u solve it on a paper then just tell me what u did (explanation) and the answers i mean summarized solution

OpenStudy (anonymous):

or just give the answers and tell me the way u did it

OpenStudy (ikram002p):

ok

OpenStudy (ikram002p):

ok i got b=3,-2 and a=2,-3 did u got the same answer ?

OpenStudy (ikram002p):

f(1)=a+(1-b)^2=6 f'(1)=2a+2(1-b)=0

OpenStudy (ikram002p):

f''(x)=2a+2

OpenStudy (anonymous):

yup i got the same thing here :) thanks alot

OpenStudy (anonymous):

but why did u derive f(x) to the second derivative ?

OpenStudy (anonymous):

but when I want to find if the critical point maximum or minimum I have to find the first derivative and I have two values for a and b so there will be two functions , right ?

OpenStudy (ikram002p):

second derivative show u if its max or min if f''(1)>0 min if f''(1)<0 max if f''(1)=0 faild so i guess its depand u say when b=3,a=2 then f(1)=2*2+2=6>0 min when b=-2,a=-3 then f(1)=2*-3+2=-2<0 max

OpenStudy (ikram002p):

sry i made a typo last lines f''(1)=6 f''(1)=-2

OpenStudy (anonymous):

thanks I know that but It didn'r come to my mind lol

OpenStudy (anonymous):

thanks a lot

OpenStudy (ikram002p):

oh ic ^^

OpenStudy (ikram002p):

np, ur wlc any time :)